### v1

# Electric field tunable magnetoexcitons in Xenes-hBN-TMDC, Xenes-hBN-BP, and Xenes-hBN-TMTC heterostructures

Abstract
In this work we propose novel van der Waals (vdW) heterostructures composed of Xenes, transition metal dichalcogenides (TMDCs), phosphorene, and transition metal trichalcogenides (TMTCs), which are separated by insulating hexagonal boron nitride (hBN) layers. We investigate theoretically the behavior of Rydberg indirect excitons in Xenes-hBN-TMDC, Xenes-hBN-BP, and Xenes-hBN-TMTC heterostructures, subject to parallel external electric and magnetic fields that are oriented perpendicular to the layers. By incorporating both isotropic and anisotropic materials, we demonstrate that excitonic properties can be effectively tuned through the external field strengths and the heterostructure design.
Our results show that the exciton reduced mass and the binding energy increase with the electric field strength, while enhanced dielectric screening from additional hBN layers reduces the binding energy. Anisotropic materials exhibit distinct excitonic responses, including variations in diamagnetic behavior. Moreover, the diamagnetic energy contributions and coefficients decrease with stronger electric fields but increase with the number of hBN layers. Finally, we explore the potential of time-periodic electric fields with Floquet band-structure engineering. These findings provide a comprehensive framework for controlling excitonic phenomena in low-dimensional materials, enabling the design of advanced optoelectronic and quantum devices.
I Introduction
Two-dimensional layered materials have attracted persistent attention and have been extensively studied over the past two decades. Graphene has sparked significant research interest in two-dimensional (2D) layered materials such as transition metal dichalcogenides (TMDCs) [1], and group IVA materials such as silicene, germanene, and stanene, collectively known as Xenes, which consist of atoms arranged in a honeycomb lattice similar to that of graphene, but with varying degrees of buckling of two sublayers with respect to each other [2, 3]. In contrast to graphene, these materials exhibit either direct or indirect band gaps and possess remarkable physical properties as a result of their crystal symmetry and reduced dimensionality. Moreover, TMDCs and Xenes typically display in-plane structural isotropy in their electrical and optical properties.
Phosphorene, the monolayer form of black phosphorus (BP), was synthesized in 2014 [4] and offers several advantages over TMDCs and Xenes. Its most notable features include strong in-plane anisotropy, a thickness-dependent band gap, and high carrier mobility. The monolayer structure of BP appears to be composed of two distinct planes forming a puckered honeycomb structure [5, 6, 7], which gives rise to its anisotropic electronic characteristics.
A newer class of 2D materials exhibiting strong in-plane anisotropy includes the group IVB transition metal trichalcogenides (TMTCs). These materials also feature a puckered honeycomb structure and have garnered increasing interest in recent years [8]. Prototypical examples such as TiS3, ZrS3, and ZrSe3 have been successfully synthesized [9]. TMTCs are composed of atomic layers held together via weak van der Waals forces and exhibit quasi-one-dimensional behavior similar to that of phosphorene. Their unique and highly anisotropic crystal structures lead to drastically different material properties along different crystallographic directions [10, 11, 12, 13, 14, 7, 15, 16, 17, 9, 18, 19].
Due to the structural differences between the armchair (AC) and zigzag (ZZ) directions, both phosphorene and TMTCs exhibit strong in-plane anisotropy. Many of their physical properties, including the effective masses of charge carriers, vary significantly between these two directions. This anisotropic nature contrasts the in-plane isotropy observed in TMDCs and Xenes. TiSe3, ZrS3, and ZrSe3 monolayers are all indirect-gap semiconductors [20, 21], whereas phosphorene is a direct-gap semiconductor with a highly anisotropic dispersion near the band edge.
The concept of exciton formation in systems with spatially separated electrons and holes—specifically in coupled quantum wells where electrons and holes occupy different wells—was first introduced in Ref. [22], inspiring a wealth of theoretical and experimental studies on double quantum wells and heterostructures composed of two 2D materials. Today, heterostructures made from layers of TMDCs, phosphorene, Xenes, TMTCs, and others are at the forefront of nanomaterials research. The formation of excitonic complexes such as excitons, trions, and biexcitons in these systems leads to unique electronic, optical, and mechanical properties, which differ significantly from their bulk counterparts and hold great promise for applications in nanoelectronics, optoelectronics, and quantum computing [23, 24].
Heterostructures comprising 2D monolayers separated by dielectric layers—particularly when there is a lattice mismatch between the constituent layers, as, for example, in MoS2–hBN–WSe2 or MoSe2–hBN–WSe2—have been extensively studied over the past two decades. These investigations have uncovered rich physics arising from interlayer coupling, moiré superlattices, and excitonic phenomena, providing deep insights into their electronic properties and excitonic behaviors, see [25, 26, 27] and references herein. Such studies underscore the critical role of lattice mismatch in designing and understanding the properties of 2D heterostructures, paving the way for novel applications in optoelectronics and quantum devices. Combining different 2D materials into heterostructures enables the tuning of electronic and optical properties via a precise control over stacking sequences, twist angles, and interlayer interactions. For example, a relative twist between layers can lead to the formation of moiré patterns that significantly influence the electronic band structure and may give rise to phenomena such as superconductivity [28] and flat electronic bands [29].
The band structures of TMDC, phosphorene, and TMTC monolayers are largely unaffected by a perpendicular electric field, and the same holds true for heterostructures composed of these monolayers. In contrast, Xenes exhibit a tunable band gap due to their buckled structure. Therefore, heterostructures such as Xenes-hBN-TMDC, Xenes-hBN-BP, or Xenes-hBN-TMTC can exhibit a tunable gap, which in turn affects the properties of indirect excitons formed within them. The buckled geometry of Xenes leads to a potential difference between sublattices, allowing control over the band gap, effective masses of charge carriers, binding energies, and diamagnetic coefficients (DMCs) of magnetoexcitons. We do not consider systems where Xene monolayers are placed directly on top other layers without hBN separation, as Xenes are unstable in air [30, 31].
In 2D systems, the reduced dielectric screening and strong Coulomb interactions enhance the binding of excitons, making them especially sensitive to magnetic fields. When a magnetic field is applied, it affects both the binding energy and the diamagnetic shift of the exciton. In van der Waals heterostructures, composed of novel 2D materials, the impact of a magnetic field on excitons is further influenced by the nature of interlayer coupling, effective masses, dielectric environments. These systems allow for tunable magnetoexcitonic behavior through effective mass anisotropy [32, 8], layer engineering, electric field control.
In this article, we propose van der Waals (vdW) heterostructures engineered by combining layers of Xenes and TMDCs, as well as Xenes and phosphorene or TMTCs, separated by hBN layers. We investigate the properties of excitons in these heterostructures under the influence of external electric and magnetic fields. The electric and magnetic fields are parallel and oriented perpendicular to the plane of the vdW heterostructure.
The goals of this article are twofold: (i) to demonstrate the tunability of the properties of indirect excitons in Rydberg states within isotropic (Xenes-hBN-TMDC) and anisotropic (Xenes-hBN-BP or Xenes-hBN-TMTC) heterostructures using external electric and magnetic fields; and (ii) to explore the use of a time-periodic electric field as a tool for band-structure engineering. The remainder of this article is organized as follows. In Section II, we present the theoretical model used to describe the electron-hole system in the presence of external electric and magnetic fields, both with and without effective mass anisotropy of the charge carriers. Section III contains the results of our calculations for the binding energies (BEs) of Rydberg exciton states under an external electric field in vdW heterostructures. In the same section, we also present and discuss the BEs and diamagnetic coefficients (DMCs) of magnetoexcitons in these heterostructures under the influence of an external electric field. Additionally, we explore a band-structure engineering via a time-periodic electric field. Finally, our concluding remarks are given in Section IV.
Figure 1: (Color online) Schematic illustration of excitons in ($a$) Xenes-hBN-TMDC and ($b$) Xenes-hBN-BP van der Waals heterostructure in electric and magnetic fields. By replacing phosphorene by TMTC monolayer one gets Xenes-hBN-TMTC heterostructure.
II Theoretical Model
It is known that electrostatically bound electrons and holes in an external magnetic field form magnetoexcitons. In this section, following Refs. [33, 34, 35, 36], we briefly introduce the theoretical model for describing Mott-Wannier magnetoexcitons in Xenes-hBN-TMDC and Xenes-hBN-BP or Xenes-hBN-TMTC heterostructures. We consider the energy contribution, $\Delta E$, from external electric and magnetic fields to the binding energies, $E_{B}$, of Rydberg states of magnetoexcitons and their diamagnetic coefficients (DMCs). In the system under consideration, excitons are confined within a van der Waals heterostructure, where $N$ layers of hBN monolayers separate Xenes and TMDC or Xenes and phosphorene or Xenes and TMTC monolayers, respectively.
Let us outline the low-energy model that describes exciton states in Xenes-hBN-TMDC and Xenes-hBN-BP van der Waals heterostructures in the $x-y$ plane. We consider an electron and a hole in parallel magnetic $\mathbf{B}=(0,0,B_{z})\equiv(0,0,B)$ and electric $\mathbf{E}=(0,0,E_{z})\equiv(0,0,E)$ fields, which are perpendicular to the heterostructure, as depicted in Fig. 1. Monolayers of silicene, germanene, and low-buckled stanene can be pictured as honeycomb graphene monolayers with an out-of-plane buckling, such that the A and B triangular sublattices are offset with respect to the $x-y$ plane with a typical distance $d_{0}$ to the latter. This distance is known as the buckling constant or buckling factor. The intrinsic sensitivity of Xenes to a perpendicular electric field is due to the offset between the two triangular sublattices. In particular, the resulting asymmetry causes an on-site potential difference between the sublattices. In the absence of an external electric field the band structure of Xenes in the vicinity of the $K/K^{\prime}$ points resembles gapped graphene, while the application of a perpendicular electric field changes the band gap in the monolayer Xenes. hus, the single-particle dispersion of quasiparticles in a monolayer Xenes is described in the vicinity of the $K/K^{\prime}$ points by the two-dimensional (2D) Hamiltonian [37]
$H_{0}=v_{F}(\xi p_{x}{\hat{\tau}}_{x}+p_{y}{\hat{\tau}}_{y})-\xi\Delta_{so}{\hat{\sigma}}_{z}{\hat{\tau}}_{z}+\Delta_{z}{\hat{\tau}}_{z},$ (1)
where ${\hat{\sigma}}_{j}$ and ${\hat{\tau}}_{j}$ are Pauli matrices in the spin and in the pseudospin space, respectively. The pseudospin is a result of the underlying honeycomb lattice. $\xi=\pm 1$ refers to the valley index of the honeycomb lattice. Moreover, $v_{F}$ is the Fermi velocity, $\Delta_{so}$ is the spin-orbit splitting, and $\Delta_{z}=ed_{0}E$ describes the potential caused by a perpendicular electric field $E$ on the two triangular sublattices of the honeycomb lattice. Since only ${\hat{\sigma}}_{z}$ appears, the Hamiltonian $H_{0}$ separates in the spin space into two massive 2D Dirac Hamiltonian with dispersions
$E_{\sigma}(p)=\pm\sqrt{\Delta_{\xi\sigma}^{2}+v_{F}^{2}p^{2}}$ (2)
with $\Delta_{\xi\sigma}=|\xi\sigma\Delta_{so}-ed_{0}E|$, where $\xi,\sigma=\pm 1$ are the valley and spin indices, respectively. Thus, we have four bands with two different gaps $\Delta_{\xi\sigma}$ due to $\xi\sigma=\pm 1$. The shift of $\Delta_{\xi\sigma}$ by the electric field $E$ was previously discussed in Refs. [38, 39, 40, 41], as well as in Ref. [37]. For the following we assume that the kinetic energy $v_{F}^{2}p^{2}$ is small compared to $\Delta_{\xi\sigma}^{2}$, such that we can expand the square root to obtain a parabolic dispersion $E_{\sigma}(p)\sim\pm(|\Delta_{\xi\sigma}|+p^{2}/2m)$ with the effective electron mass $m=|\Delta_{\xi\sigma}|/v_{F}^{2}$ [42, 2]. Thus, the conduction and valence bands are parabolic in the vicinity of the $K/K^{\prime}$ points. The effective masses of electrons and holes are the same due to the symmetry between the lowest conduction and highest valence bands.
At non-zero electric fields, both the valence and conduction bands, split into upper bands with a large gap (when $\xi=-\sigma$), and lower bands with a small gap (when $\xi=\sigma$). We call the excitons formed by charge carriers from the large gap $A$ excitons, and those formed by charge carriers in the small gap $B$ excitons. At small electric fields, germanene and especially stanene show significant differences between the masses of the $A$ and $B$ excitons. The latter is due to their large intrinsic band gaps. Silicene, which has an intrinsic band gap on the order of a couple of meV, exhibits very little difference between the masses of $A$ and $B$ excitons, even at relatively small electric fields. At large electric field the difference between the $A$ and $B$ exciton mass is negligible in silicene and germanene. In all cases, the mass of the $A$ exciton exceeds the mass of the $B$ exciton. TMDC monolayers are isotropic and effective masses of an electron and hole are different and $m_{h}>m_{e}$.
In the next step we will consider a system whose translation invariance is broken by an external magnetic field. To this end, let us introduce the coordinate vectors of the electron and hole for the Mott-Wannier exciton in the Xenes and TMDC (phosphorene or TMTC) layers. The following in-plane coordinates $\mathbf{r}_{e}(x_{e},y_{e})$ for an electron in Xenes layer and $\mathbf{r}_{h}(x_{h},y_{h})$ for a hole in TMDC or phosphorene layer, respectively, are used in our description. The electron/hole spins are polarized by the magnetic field and the honeycomb structure is irrelevant because the system is dominated by the magnetic length. Then the Hamiltonian of an interacting pair of an electron at the site $r_{e}$ and a hole at the site $r_{h}$ reads ($\hbar=c=1$)
$\hat{H}=$ $\frac{1}{2m^{e}}\Big{(}i\nabla_{x}^{e}-eA_{x}(r_{e})\Big{)}^{2}+\frac{1}{2m^{e}}\Big{(}i\nabla_{y}^{e}-eA_{y}(r_{e})\Big{)}^{2}+\frac{1}{2m_{x}^{h}}\Big{(}i\nabla_{x}^{h}+eA_{x}(r_{h})\Big{)}^{2}+\frac{1}{2m_{y}^{h}}\Big{(}i\nabla_{y}^{h}+eA_{y}(r_{h})\Big{)}^{2}$ (3)
$+V\Big{(}|\mathbf{r}_{e}-\mathbf{r}_{h}|\Big{)},$
where $V\left(|\mathbf{r}_{e}-\mathbf{r_{h}}|\right)$ describes the Coulomb interaction between the electron and hole via a cental Coulomb potential
$V\left(\sqrt{x^{2}+y^{2}+D^{2}}\right)=-\frac{ke^{2}}{\kappa\left(\sqrt{x^{2}+y^{2}+D^{2}}\right)},$ (4)
where $r^{2}=x^{2}+y^{2}$. The effective masses $m^{e}$ and $m^{h}$ are taken from the calculation of the expanded square-root dispersion of Eq. (2). The anisotropic nature of the phosphorene and TMTC, in contrast to other 2D isotropic materials such as graphene and TMDC semiconductors, breaks the circular symmetry. This requires the use of Cartesian coordinates for the description of excitons. The asymmetry of the electron and hole dispersion in phosphorene is reflected in the Hamiltonian for the Mott-Wannier exciton, and the Hamiltonian for an interacting electron-hole pair in Xenes-hBN-BP or Xenes-hBN-TMTC heterostructures in the external electric and magnetic fields is obtained within the framework of the effective mass approximation. In particular, the electric field dependence of Hamiltonian (3) is implicit via the effective mass of electron. Equation  (4) describes the interaction between the electron and the hole that are located in different Xenes and phosphorene or TMTC monolayers, separated by a distance $D=h+Nl_{\text{hBN}}$. $l_{\text{hBN}}=0.333$ nm is the thickness of the hBN layer and $h$ is the thickness of the phosphorene or TMTC, while $m_{x}^{h}$ and $m_{y}^{h}$ correspond to the effective mass of the hole in the $x$ or $y$ direction in the phosphorene or TMTC, respectively. In the case of Xenes TMDC heterostructure $m_{x}^{e}=m_{x}^{h}=m_{e}$ and the Hamiltonian (3) becomes simpler.
Following the standard procedure [43] for the separation of the relative motion of the electron-hole pair from their center-of-mass motion, one introduces variables for the center-of-mass of an electron-hole pair. After a lengthy calculation, following Refs. [44, 34, 35, 36], one obtains the equation that describes the Mott–Wannier exciton in the Rydberg optical states in the external electric and magnetic field perpendicular to the Xenes-hBN-BP heterostructure. Finally, the equation for the relative motion of the electron and hole in the Xenes-hBN-BP heterostructure with zero center-of-mass momentum [33, 44, 45] reads
$\left[-\frac{1}{2\mu_{x}}\frac{\partial^{2}}{\partial x^{2}}-\frac{1}{2\mu_{y}}\frac{\partial^{2}}{\partial y^{2}}+\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}+\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}+V(x,y)\right]\Phi(x,y)=\mathcal{E}\Phi(x,y).$ (5)
where $\mu_{x}=\frac{m^{e}m_{x}^{h}}{m^{e}+m_{x}^{h}}$ and $\mu_{y}=\frac{m^{e}m_{y}^{h}}{m^{e}+m_{y}^{h}}$ are the reduced masses in the $x$ and $y$ directions, respectively. In Eq. (5) the anisotropy is present in the kinetic and magnetic terms, while the potential term has isotropic form and the action of the electric field is present via anisotropic reduced masses $\mu_{x}$ and $\mu_{y}$. We solve Eq. (5) using $V\left(\sqrt{x^{2}+y^{2}+D^{2}}\right)$ for indirect magnetoexcitons. For indirect magnetoexcitons, both $V_{RK}$ and $V_{C}$ potentials are used to investigate the importance of using correct potential describing electron-hole interactions. Note that Eq. (5) does not explicitly contain any spin- or valley-dependent Zeeman terms.
The Schrödinger equation with Hamiltonian (3) for Xenes-hBN-TMDC heterostructure when $m_{x}^{e}=m_{x}^{h}=m_{h}$ has the form: $\hat{H}\Psi(\mathbf{r}_{e},\mathbf{r}_{h})=E\Psi(\mathbf{r}_{e},\mathbf{r}_{h})$, where $\Psi(\mathbf{r}_{e},\mathbf{r}_{h})$ and $E$ are the eigenfunction and eigenenergy, respectively. Due to the spherical symmetry, in this case the corresponding equation easier write in spherical coordinates. Finally, after separating the angular variable the equation with zero center-of-mass momentum reads [44, 45]:
$\left[-\frac{1}{2\mu}\frac{{\partial}^{2}}{\partial r^{2}}-\frac{1}{2\mu}\frac{1}{r}\frac{{\partial}}{\partial r}+\frac{e^{2}}{8\mu}B^{2}r^{2}+V(r)\right]\Phi(r)=E\Phi(r).$ (6)
where $\mu=\frac{m_{e}m_{h}}{m_{e}+m_{h}}$ is the reduced mass. Equation (5) describes the Mott–Wannier magnetoexciton in Rydberg optical states in 2D materials. This equation has a long history of the solution in the case of the electron-hole Coulomb interaction [46, 33, 47, 48, 49, 50, 51, 44, 52, 45].
In addition to the exciton binding energy, another important quantity is the diamagnetic coefficient, which gives the change of the exciton binding energy by an external magnetic field:
$\Delta E=|E(B)-E_{0}|,$ (7)
where $E(B)$ is the binding energy of magnetoexciton, and $E_{0}$ is the binding energy of a conventional exciton. The Taylor expansion of $\Delta E$ in powers of $B$ can be used to determine the diamagnetic coefficient. Considering the first three terms of the Taylor series [53, 54, 55] we get
$E(B)=E_{0}+\gamma_{1}B+\gamma_{2}B^{2}+O(B^{3}),$ (8)
where the second and third term are the Zeeman and diamagnetic shifts, respectively. For Rydberg states our model has only contribution from the diamagnetic shift. The diamagnetic coefficient is defined as [56, 57, 53, 58, 59, 60, 61]:
$\gamma_{2}=\frac{e^{2}}{8\mu}\langle r^{2}\rangle,$ (9)
where $\langle r^{2}\rangle$ is the expectation value of $r^{2}$ over exciton envelope wave function. The diamagnetic coefficient quantifies the extent to which a material resists changes in its magnetization when subjected to an external magnetic field. The value of a diamagnetic coefficient depends on the exciton binding energy, because the strength of the exciton binding energy affects how it responds to an external field. The magnetic field strength alters the size of the exciton wavefunction and modifies the energy levels. The dimensionality and the structure of the material affect the exciton radius and lead to a different diamagnetic response.
Figure 2: (Color online) The dependence of the reduced mass of an exciton in Xenes/TMDC ($a$) and Xenes/phosphorene ($b$) heterostructures on the electric field.
[…]
III Results of Calculations and Discussion
First, let us focus on the reduced mass in the Xenes-hBN-TMDC and Xenes-hBN-BP heterostructures. The dependence of the exciton reduced mass in these heterostructures is depicted in Fig. 2. Analysis of the results in Fig. 2 reveals a distinct behavior of the reduced masses. In the Xenes-hBN-TMDC heterostructure, the exciton reduced mass is smaller than that in the WS2 for electric fields $E_{\perp}<1.5$ V/m, but becomes larger for $E{\perp}>1.5$ V/m. In contrast, the reduced masses $\mu_{x}$ and $\mu_{y}$ in the Xenes-hBN-BP heterostructure remain smaller than those in phosphorene across the entire range of electric field values.
[…]
The relationship between BE and reduced mass has significant implications in monolayer physics. In TMDC-hBN-TMDC heterostructures, which exhibit lower electron and hole effective masses than Xenes-hBN-TMDC at $E_{\perp}>1.5$ V/m, excitons tend to have lower binding energies. This characteristic behavior affects the efficiency of exciton-based devices, as lower binding energies facilitate the dissociation of excitons into free charge carriers—an essential process for efficient charge transport. Understanding this dependence is crucial for the design of materials with tailored electronic and optical properties, particularly in systems where excitonic effects play a dominant role.
Figure 4: (Color online) The dependence of the total potentials $W_{1}(x,y)=\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}+\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}+V(x,y)$ and $W_{2}(x,y)=\frac{e^{2}}{8\mu}B^{2}(x^{2}+y^{2})+V(x,y)$ acting on the electron-hole pair in Xenes-hBN-phosphorene and Xenes-hBN-TMDC ($a$) and Xenes-hBN-TMTC and Xenes-hBN-TMDC ($b$) heterostructures. Calculations are performed for the electron-hole pair in the external magnetic field $B=30$ T and electric field $E_{\perp}=0.5$ V/m.
Figure 5: (Color online) Dependence of energy $\Delta E$ for magnetoexcitons in $1s$ and $3s$ states in Xenes/MoSe2 and Xenes-hBN-BP heterostructures on the electric and magnetic fields.
In the next step, let us consider indirect magnetoexcitons formed by electron-hole pair in van der Waals heterostructures (e.g., Xenes–hBN–TMDC, Xenes–hBN–BP, Xenes–hBN–TMTC), in the presence of magnetic field. In these vdW heterostructures magnetoexcitons emerge as a result of the influence of an external magnetic field on excitons, i.e. the presence of an external magnetic field alters the excitonic states and leads to the formation of magnetoexcitons, which are excitons whose properties are modified by the magnetic field. The characteristics of magnetoexcitons are governed by both the underlying electronic structure of the heterostructure and the strength of the magnetic field, modifying their internal structure and energy spectrum. In Eqs. (5) and (6) $W_{1}(x,y)=\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}+\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}+V(x,y)$ and $W_{2}(x,y)=\frac{e^{2}}{8\mu}B^{2}(x^{2}+y^{2})+V(x,y)$ are total potentials that act on the electron-hole pair in these heterostructures, respectively. The Coulomb interaction $V(x,y)$ (3) as well as $W_{2}(x,y)$ have a rotational symmetry. In contrast, though, the terms $\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}$ and $\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}$ in $W_{1}(x,y)$ break the rotational symmetry due to its anisotropic nature. As a result of reduced masses anisotropy, when $\mu_{x}>\mu_{y}$, $\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}$ gives larger contribution due to the magnetic field than the term $\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}$ and vice versa if $\mu_{y}>\mu_{x}$. Thus, the electron and hole masses anisotropy leads to the anisotropy of $W_{1}(x,y)$. The dependence of the total potential $W_{1}(x,y)$ and $W_{2}(x,y)$, respectively, acting on the electron-hole pair on $x$ and $y$ coordinates is shown in Fig. 4. Let us emphasize that the anisotropy of the potential $W_{1}(x,y)$ with respect to $W_{2}(x,y)$ is different in the case of heterostructures with phosphorene and TMTC.
The results of calculations of the energy contribution $\Delta E$ due to a magnetic field at different values of the electric field are presented in Fig. 5. Here, we consider MoSe2 and phosphorene as representative cases for isotropic and anisotropic materials, respectively. Contributions of the magnetic field are proportional to the factors $1/\mu$ for an isotropic material and $1/\mu_{x}$ and $1/\mu_{y}$ for an anisotropic material. Therefore, the electric field increase leads to the decrease of $1/\mu$, $1/\mu_{x}$, and $1/\mu_{y}$, consequently, $\Delta E$ decreases. At the small electric field $1/\mu$, $1/\mu_{x}$, and $1/\mu_{y}$ are comparable, although $1/\mu_{x}$, and $1/\mu_{y}$ a little bit exceed $1/\mu$. By increasing the number of hBN layers, the electrostatic attraction decreases, which changes $\Delta E$. When $N=1$ the difference between $\Delta E$ for both heterostructures is negligible (Fig. 5 $a$ $b$). When $E_{\perp}>1$ V/m, $1/\mu<1/\mu_{x}$ and $1/\mu<1/\mu_{y}$, and $\Delta E$ becomes larger for heterostructure with phosphorene (Fig. 5 $c$ $f$). In Fig. 5, we demonstrate this for $1s$ and $3s$ states, but the same tendencies can be extended to $2s$ and $4s$ states. In Xenes–hBN–MoSe2 and Xenes–hBN–BP heterostructures, the strong field-tunable band gaps and anisotropic effective masses of phosphorene can lead to pronounced shifts in exciton binding energy and modified magnetoexciton behavior compared to conventional heterostructures build with TMDC or phosphorene monolayers. Due to the larger effective dielectric constant introduced by hBN layers, excitons are more weakly bound, making their magneto-optical response more sensitive to external fields.
The analysis of these calculations leads to the following conclusions: (i) the energy contribution increases due to the increase of magnetic field; (ii) The addition of hBN layers gives an increase in $\Delta E$ energy; (iii) The increase in the electric field leads to the decrease of $\Delta E$.
The diamagnetic coefficients for excitons in Si-hBN-MoSe2 and Si-hBN-BP heterostructures, obtained in the framework of our approach, are reported in Table 1. For Xenes-hBN-TMDC we show a representative case - Si-hBN-MoSe2. Analysis of the results presented shows: i. the diamagnetic coefficients are decreasing with increasing electric field for all Rydberg states in both heterostructures; ii. the increase in the number of insulating hBN layers leads to the increase of $\sigma$ for $1s$, $2s$ and $3s$ states; iii. $3s$ and $4s$ states generally exhibit high diamagnetic coefficients due to the increase exciton’s $r^{2}$ affected by the size of the exciton wavefunction.
Table 1: Diamagnetic coefficients for magnetoexcitons in Si-hBN-phosphorene and Si-hBN-MoSe2. DMCs are obtained when $R^{2}=0.9998$ for the linear regression model.
Diamagnetic coefficient $\gamma_{2}$ [$\mu$eV/T2]
Si/hNB/BP Si/hBN/MoSe2
$N=1$
$E_{\perp}$ [V/m] 0.1 1.0 2.0 0.1 1.0 2.0
[…]
$N=6$
[…]
Our results suggest the consideration of a time-periodic electric field with period $T$ that, according to our single-exciton calculation, creates a time-periodic exciton mass. Then the resulting time-periodic Hamiltonian $H_{\bf p}(t)$ can be treated within the Floquet approach [62]. This yields a Floquet Hamiltonian, which describes a single exciton. However, a more realistic situation is to consider a condensate of excitons. Having calculated the properties of a single exciton in external fields, we can extend this case to an excitonic condensate in the double-layer structure. Assuming that the excitons have a small dipolar moment and interact weakly with each other, we can ignore the exciton-exciton interaction in the first order approximation and calculate the band structure of the non-interacting excitonic gas. To maintain translational invariance, we simplify the situation further and consider the case without an external magnetic field. This leads to a translation-invariant distribution of excitons, which we describe in Fourier representation with momentum ${\bf p}$ as the effective Hamiltonian [22], which was mentioned at the beginning of Sect. II:
$H_{\bf p}(t)=\begin{pmatrix}p^{2}/2\mu(t)&\Delta_{p}\\ \Delta_{p}&-p^{2}/2\mu(t)\end{pmatrix},$ (10)
where $\Delta_{p}$ represents the uniform band splitting due to electron-hole pairing and $\mu(t)$ the time-dependent reduced mass. $\Delta_{p}$ can be calculated self-consistently, following the approach of Ref. [22], where the mass oscillations are corrections in comparison to the time-average mass. At low density, $\Delta_{p}>0$ for a sufficiently strong interlayer attraction between electrons and holes. The other condition is that the gap $2\Delta_{p\xi\sigma}=2|\xi\sigma\Delta_{so}-ed_{0}E_{0}\cos wt|$ must be larger than the kinetic energy $p^{2}/2\mu$. This means that $E_{0}$ must be smaller than $\Delta_{so}/ed_{0}$. Thus, we can assume within the leading approximation that only the reduced mass $\mu(t)$ depends on the external electric field. From $H_{\bf p}(t)$ we obtain the Floquet Hamiltonian [62] as
$H_{F;{\bf p}}=\frac{i}{T}\log\left[\prod_{t=0}^{T}e^{-i{\tilde{H}}_{\bf p}(t)dt}\right].$ (11)
Its eigenvalue spectrum yields the effective band structure of the Floquet approach. This can be understood in the present case as if the variation of the reduced mass $\mu(t)$ increases and decreases the bandwidth periodically in time. Since the Floquet Hamiltonian is created by the unitary operator, values of the spectrum of $H_{\bf p}(t)$ that are outside the interval $[-\pi,\pi]$ are backfolded on the complex unit circle. This leads to the restriction of the bandwidth to $2\pi$. This effect is illustrated in Fig. 6, which reveals that the time-dependent electric field applied to the layered system provides a tool for band-structure engineering, similar to the concept presented in Ref. [63].
Figure 6: (Color online) Typical effect of Floquet band engineering: while ($a$) represents the static case, ($b$) is the result of the Floquet dynamics due to a periodically time-dependent exciton mass, which represents a backfolded spectrum on the interval $[-\pi,\pi]$.
IV Concluding remarks
In this work we suggest the novel vdW heterostructures, comprising Xenes, TMDCs, phosphorene, and TMTCs monolayers separated by hBN insulating layers. We have investigated theoretically the properties of Rydberg indirect excitons in Xenes-hBN-TMDC, Xenes-hBN-BP, and Xenes-hBN-TMTC heterostructures subjected to parallel external electric and magnetic fields that are oriented perpendicular to the layers. By incorporating both isotropic and anisotropic materials, we demonstrated the ability to control and tune excitonic properties through external fields and structural parameters.
Our study highlights several key findings: (i) the exciton reduced mass and, consequently, the binding energy can be effectively modulated by the strength of the electric field and the number of hBN layers; (ii) anisotropy in the effective mass leads to distinguishable trends in binding energy and diamagnetic response when compared to isotropic systems; and (iii) the energy contribution from the magnetic field and the diamagnetic coefficients decrease with increasing electric field, while increasing the number of hBN layers enhances these quantities due to reduced Coulomb interaction.
Furthermore, we explored the application of a time-periodic electric field as a potential method for band structure engineering. This opens avenues for dynamic control of excitonic states in 2D material systems, paving the way for novel optoelectronic and quantum devices based on excitonic phenomena.
In summary, vdW heterostructures, composed of two layers of 2D materials when one of the layers is Xenes, offer a versatile platform for exploring novel electronic and optical properties. By carefully engineering the stacking, orientation, and strain of these materials and applying external electric and magnetic fields, researchers can design systems with tailored properties, which are suitable for a wide range of applications in next-generation electronic and optoelectronic devices.